# What Variables Affect Out-of-Distribution Generalization in Pretrained Models?

### v2

@@ C.11 Large Pre-trained Models Do Not Exhibit The Tunnel Effect
Figure 16: OOD performance of large pre-trained models. We show normalized linear probe accuracy on the ID (ImageNet-1K) and OOD datasets. The OOD curve is the average of 8 OOD datasets. Shaded area denotes standard deviation ($\pm$). Most models do not show any tunnel effect.
@@
Table 8: Final OOD Accuracy of Large Pre-trained Models. We report average results with 95% confidence intervals (CI). ImageNet-R is abbreviated as IN-R. $\gamma$ denotes over-parameterization level.
@@
In Fig. 16 and Table 9, we observe that the tunnel is not present in most cases except for the ResNet-50 models. Surprisingly, we do not see much difference between SL and SSL models in terms of % OOD performance retained. This suggests that the tunnel effect is mitigated by variables shared among SL and SSL such as augmentations, image resolution, and ID class count. As shown in Table 8, the SSL method, MUGS achieves the highest average OOD accuracy among all methods. In terms of DNN architecture, ViT-B exhibits better OOD performance than ResNet-50. Interestingly, among CNNs, ConvNeXt-B, which adapts design choices from Swin-transformer [87], shows better OOD performance than ResNet-50. They also rival ViTs. This corroborates our argument that DNN architecture matters for OOD performance.
@@
The comparison between SSL and SL models is not straightforward as there are multiple factors involved, e.g., pretext tasks, training epochs, augmentations, and so on. Except for MAE, other SSL methods exhibit strong OOD performance compared to SL counterparts (Table 8). Previous works [52, 88] also found that MAE performs poorly in linear evaluation. When we compared SL vs. SSL models in the same family, in terms of average OOD accuracy, SSL ($71.42\%$) outperforms SL ($68.83\%$). We report the ID accuracy of these models in Table 10. Given that SSL methods employ more augmentations than their SL counterparts (see Table 3), the performance improvements of SSL methods over SL ones are likely, in part, due to stronger augmentations, which is also argued in prior studies [89, 90]. While complimenting earlier works, our work confirms that large pre-trained models do not form tunnels, adding valuable perspectives to current knowledge.
C.12 High Resolution Does Not Exhibit Compression
Figure 18: Compression analysis. The t-SNE comparison between VGGm-11 models trained on low- (1st row) and high-resolution (2nd row) images of the same ID dataset (ImageNet-100) in an augmentation-free setting. Layer 8 marks the start of the tunnel in VGGm-11 trained on $32\times 32$ images whereas $224\times 224$ resolution does not create any tunnel. Layer 10 is the penultimate layer. The tunnel layers (layer 8-10) progressively compress representations for $\mathbf{32\times 32}$ resolution whereas corresponding layers for $\mathbf{224\times 224}$ resolution do not exhibit similar compression. For clarity, we show 5 classes from ImageNet-100 and indicate each class by a distinct color. The formation of distinct clusters in the $32\times 32$ model is indicative of representation compression and intermediate neural collapse [10], which impairs OOD generalization.
In this section, we compare the t-SNE results of the VGGm-11 models trained on ImageNet-100 with $32\times 32$ and $224\times 224$ input resolutions in an augmentation-free setting. For comparison, we select five classes from the ImageNet-100 dataset. We extract embeddings from tunnel layers to analyze compression. We illustrate these embeddings using t-SNE plots. Fig. 18 demonstrates that a model trained on higher resolution images exhibits weaker compression as opposed to a model trained on lower resolution images. Thus, the characteristic of the tunnel layers i.e., feature compression emerges for low-resolution inputs only.
C.13 More Resolution Results
@@
Here we present more resolution results. The image resolution shows a significant impact on the tunnel effect and OOD transferability.
@@
Appendix D Classes of ImageNet-100 ID Dataset
Appendix E In-Distribution Performance
Table 11 has the results for each of the 64 DNN models studied in our main results. For each model and ID dataset combination, it gives the ID accuracy for the resolution/augmentation combination.
[…]
@@ Appendix E In-Distribution Performance
Table  10 shows the ID performance of 8 ImageNet-1K pre-tained models.

### v3

@@ C.11 Large Pre-trained Models Do Not Exhibit The Tunnel Effect
Figure 17: OOD performance of large pre-trained models. We show normalized linear probe accuracy on the ID (ImageNet-1K) and OOD datasets. The OOD curve is the average of 8 OOD datasets. Shaded area denotes the standard deviation. Most models do not show any tunnel effect.
@@
Table 8: OOD Accuracy of Large Pre-trained Models. We report average results with 95% confidence intervals (CI). ImageNet-R is abbreviated as IN-R. $\gamma$ denotes over-parameterization level.
@@
In Fig. 17 and Table 9, we observe that the tunnel is not present in most cases except for the ResNet-50 models. As shown in Table 8, the SSL method, MUGS achieves the highest average OOD accuracy among all methods. In terms of DNN architecture, ViT-B exhibits better OOD performance than ResNet-50. Interestingly, among CNNs, ConvNeXt-B, which adapts design choices from Swin-transformer [95], shows better OOD performance than ResNet-50. They also rival ViTs. This corroborates our argument that DNN architecture matters for OOD performance.
@@
The comparison between SSL and SL models is not straightforward as there are multiple factors involved, e.g., pretext tasks, training epochs, augmentations, and so on. Except for MAE, other SSL methods exhibit strong OOD performance compared to SL counterparts (Table 8). Previous works [52, 96] also found that MAE performs poorly in linear evaluation. When we compared SL vs. SSL models in the same family, in terms of average OOD accuracy, SSL ($71.42\%$) outperforms SL ($68.83\%$). As illustrated in Fig. 18, SSL models perform better than SL models in terms of % OOD performance retained. We report the ID accuracy of these models in Table 10. Given that SSL methods employ more augmentations than their SL counterparts (see Table 3), the performance improvements of SSL methods over SL ones are likely, in part, due to stronger augmentations, which is also argued in prior studies [97, 98]. While complimenting earlier works, our work confirms that large pre-trained models do not form tunnels, adding valuable perspectives to current knowledge.
C.12 Examining Actual Accuracy
(a) Resolution
(b) Augmentation
(c) ID Class Count
Figure 19: Dataset variables. Impact of resolution, augmentation, and ID class count on ID accuracy, OOD accuracy, and ID/OOD alignment. This analysis is based on actual accuracy (no normalization) averaged over models and datasets. The accuracy and alignment score range from 0 to 1.
(a) Depth & Overparam. Level
(b) Spatial Reduction
(c) Stem
Figure 20: DNN architecture variables. Impact of depth (D), overparameterization level ($\gamma$), spatial reduction ($\phi$), and stem on ID accuracy, OOD accuracy, and ID/OOD alignment. This analysis is based on actual accuracy (no normalization) averaged over models and datasets. The accuracy and alignment score range from 0 to 1.
We also examine the impact of different variables and interventions on the ID performance, OOD performance, and ID/OOD alignment by analyzing the actual accuracy (no normalization). We show these results in Fig. 19 and Fig. 20. We observe that the actual accuracy trends are consistent with our previous findings where variables that improve (degrade) OOD generalization also show higher (lower) ID/OOD alignment.
C.13 Additional Resolution Results
@@
Here we present additional resolution results. The image resolution shows a significant impact on the tunnel effect and OOD transferability. As shown in Fig. 19(a), higher resolution images achieve higher performance in terms of ID accuracy, OOD accuracy, and ID/OOD alignment.
@@
Rank Analysis. We examine the numerical rank of the representations evaluated on the ID dataset, following prior work [9]. In Fig. 25, we observe that a model trained on high-resolution images maintains a higher rank than a model trained on low-resolution images, corroborating our previous findings (Sec. 4.1.3). The representations rank for resolution $32\times 32$ plummets after the extractor, exhibiting the neural collapse phenomenon [10] whereas resolution $224\times 224$ retains a much higher rank in the corresponding layers. The degradation in OOD accuracy and ID representation rank is more pronounced for low-resolution images than for high-resolution images. As a result, a model trained on low-resolution images exhibits a stronger tunnel effect than one trained on high-resolution images.
(a) Resolution $32\times 32$ (strong tunnel effect)
(b) Resolution $224\times 224$ (weak tunnel effect)
Figure 25: High-resolution model maintains a high representation rank. This figure corresponds to Fig. 1 in the main text. In this setting, identical VGGm-17 architectures are trained on identical ID datasets (100 ImageNet classes), where only the resolution is changed. The Y-axis shows the actual accuracy (no normalization) of linear probes trained on ID and OOD datasets. The OOD curve is the average of 8 OOD datasets (Sec. 3.3), with the standard deviation denoted with shading. The Y-axis also shows the numerical rank evaluated on the ID dataset. Here the OOD accuracy and ID representation rank align with the dynamics of the tunnel effect: the stronger the tunnel effect, the lower the OOD accuracy and ID representation rank.
Appendix D In-Distribution Performance
In this section, we present the ID performance of all pre-trained DNNs. Table  10 shows the ID performance of 8 large-scale off-the-shelf models pre-tained on ImageNet-1K dataset.
@@ Appendix D In-Distribution Performance
Table 11 presents the results for each of the 64 DNN models studied in our main results. For each model and ID dataset combination, it gives the ID accuracy for the resolution/augmentation combination.
[…]
@@
Appendix E Additional Statistical Results
In this section, we report additional statistical results for different variables and interventions. Table 12 includes average results with confidence interval. Table 13 reports effect size (Cliff’s Delta) and p-value (Wilcoxon signed-rank test) for various pair comparisons where #$\#E$ denotes the number of paired experiments (total number of experiments is $2\times E$).